My Friend and I have tried to find the inverse of this problem but we can't seem to get anything to work. our teacher hasn't discussed this with the class and now we have a project due monday on it.
the equation is BIM= W*704.5/H^2
BIM= Body Mass indicator, a mathematical formula used to determine how healthy an individual is
W= Weight of person in pounds
H= Height of person in inches
we are looking for an inverse that we have the weight but we need to find the hieght. the person weights 100 lbs.
we've tried changing the equation into (100*704.5)*H^-2 among many other things. thanks for helping if you can.

Answer by stanbon(75887)

BIM= W*704.5/H^2
You don't want the inverse of this.
You want to solve the equation for H.
1st: Interchange BIM and H^2
H^2 = 704.5*W/BIM
Take the square r=Root to get:
H=sqrt[[704.5W/BIM]
Cheers,
Stan H.